the perimeter of a rectrangular backyard is 6x+6 yards. if the width x yards, find a binomial that represents the length.
Answer by stanbon(75887)

the perimeter of a rectrangular backyard is 6x+6 yards. if the width x yards, find a binomial that represents the length.
---------------------
Perimeter = 2(Width + Length)
6x+6 = 2(x + L)
3x+3 = x + L
L = 2x+3
-----------------
Cheers,
Stan H.